Solve the proportion. Where necessary, round to the nearest hundredth.
x/4 = 6/7

Answer by TimothyLamb(4379)

x/4 = 6/7
---
x = 4*6/7
---
x = 24/7
x ~= 3.43
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php